OpenStudy (boxman61):

solve log(2x+20)=1 process please, trying to learn

OpenStudy (australopithecus):

log(x) = y is the same as x = 10^(y) thus 10^(log(2x+20))=10^(1) = 2x+20 = 10^(1) = 2x+20 = 10

OpenStudy (boxman61):

\[2x+20=10\]\[2x=-10\]\[x=-5\]

OpenStudy (boxman61):

just remembered that...thank you

OpenStudy (australopithecus):

just make both sides exponents to the power of 10 the logarithm will cancel out because it is base 10 and you can solve if you had ln(x) = y or log_8(x) = y then we would have to put e^(ln(x)) = e^(y) = x = e^y and the other example with log base 8 would be 8^y = x